Grade 12th passPhysical Chemistry

The mole fraction of solvent in 0.1 molal aqua solution

purnachandra
9 Years agoGrade 12th pass

1 Answer

Khimraj
9 Years ago
0.1 molal means 1mol solute in 10 kg water so mole of water = 10000/18 so mole fraction of solvent is (10000/18)/1 + (10000/18) = 10000/10018.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer